OpenStudy (anonymous):

How to find the volume of the solid generated by rotation about the x-axis of the part of the curve of y=sin x, between x=0 and x=pi ?

OpenStudy (kinggeorge):

\[\pi \int\limits_{0}^{\pi} \sin^2x \quad dx\]

OpenStudy (unklerhaukus):